what is the answer of these using elimination method
eq1 {3x-3000y=3}
eq2 {2x+1y=11oo
pls help me to solve

Answer by stanbon(75887)

elimination method
3x-3000y = 3
2x+ y= 1100
-----------------
Modify::
6x - 6000y = 6
6x - 3y = 3300
---------------
Subtract and solve for "y":
5997y = 3294
y = 3294/5997
------------------
y = 0.5493
----
Solve for "x" using 2x + y = 1100
2x = 1100-0.5493
2x = 1099.45
x = 549.73
-------------------
Cheers,
Stan H.
-------------------